# smallequation

### v9

@@ Abstract
Maximizing a target variable as an operational objective in a structural causal model is an important problem. Existing Causal Bayesian Optimization (CBO) methods either rely on hard interventions that alter the causal structure to maximize the reward; or introduce action nodes to endogenous variables so that the data generation mechanisms are adjusted to achieve the objective. In this paper, a novel method is introduced to learn the distribution of exogenous variables, which is typically ignored or marginalized through expectation by existing methods. Exogenous distribution learning improves the approximation accuracy of structural causal models in a surrogate model that is usually trained with limited observational data. Moreover, the learned exogenous distribution extends existing CBO to general causal schemes beyond Additive Noise Models (ANMs). The recovery of exogenous variables allows us to use a more flexible prior for noise or unobserved hidden variables. We develop a new CBO method by leveraging the learned exogenous distribution. Experiments on different datasets and applications show the benefits of our proposed method.
@@ 1 Introduction
Bayesian Optimization (BO) has broad applications, such as automating modern industrial processes, drug discovery, and more general synthetic biology, that require optimizing black-box objective functions [15; 3]. In many real-world problems, we have some structural knowledge of the unknown functions that can be leveraged to improve BO efficiency. CBO has been developed to utilize the structural information of the objective [2; 1; 23]. CBO combines ideas from causal inference, uncertainty quantification, and sequential decision-making. It generalizes Bayesian Optimization, which treats the input variables of the objective function as independent, to scenarios where causal information between input variables is available [2]. CBO has been developed for optimizing medical and ecological interventions as illustrative examples [2; 1].
@@ 1.1 Approach and Contributions
In this paper, we develop a Causal Bayesian Optimization method, EXCBO. Given observation data points of a structural causal model (SCM, [19; 18]), the exogenous variable associated with each endogenous node is recovered with an encoder-decoder framework. The learned distribution of the exogenous variable is then approximated with the density of the recovered exogenous variable that is parameterized with a flexible model, e.g. Gaussian Mixture. The recovered exogenous distribution improves the surrogate model’s accuracy in the SCM approximation. Different from existing CBO methods [2; 1; 23] that typically focus on Additive Noise Models (ANMs, [11]), our method enables us to extend CBO to general causal models beyond ANMs. Our flexible surrogate model has the benefit of better causal inference regarding CBO updating.
@@
  • We develop an EXogenous distribution learning augmented Causal Bayesian Optimization algorithm (EXCBO), where the flexibility of learning exogenous distributions naturally extends the proposed CBO framework to general causal models beyond ANMs.
@@
In the following parts of this paper, Section 2 presents the background and related work; Section 3 sets up the problem studied in this paper with the introduction of the proposed CBO framework; Section 4 presents the exogenous recovery method; The proposed SCM algorithm, EXCBO, is detailed in Section 5 with theoretical analysis of EXCBO given in Section 6; Section 7 presents experimental results; Section 8 concludes the paper.
2 Background and Related Work
@@ 2.1 Structural Causal Model
An SCM is denoted by $\mathcal{M}=(\mathcal{G},\mathbf{F},\mathbf{X},U)$. Here $\mathcal{G}$ is a directed acyclic graph (DAG), and $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$ stands for $d+1$ mechanisms of all variables. $\mathbf{X}$ is the endogenous variable set, and $U$ is the exogenous (background) variable set. Samples of the $i$th endogenous variable are generated with
@@
Here $[d]=\{0,1,...,d\}$. We use $X_{i}$ to represent the $i$th variable as well as the $i$th node in $\mathcal{G}$, and $\mathbf{pa}(i)$ stands for the parents of $i$, and $\mathbf{ch}(i)$ is the children of node $X_{i}$ regarding $\mathcal{G}$. In addition, we have $U_{i}\perp\!\!\!\perp X_{i}$ and $U_{i}\perp\!\!\!\perp U_{j},\forall i\neq j$. $f_{i}$ is a mapping from $\mathbb{R}^{|\mathbf{pa}(i)|+1}\rightarrow\mathbb{R}$. Moreover, $\mathcal{X}_{i}$ and $\mathcal{Z}_{i}$ stand for the value range of $X_{i}$ and $\mathbf{Z}_{i}$, respectively. Existing CBO methods [2; 23] usually assume that the exogenous variable is an ANM [11] to the data generation mechanism, i.e., $X_{i}=f_{i}(\mathbf{Z}_{i})+U_{i},\ \text{with}\ U_{i}\sim\mathcal{N}(0,1).$
@@ 2.2 Intervention
In SCM $\mathcal{M}$, let $\mathbf{I}$ be a subset of endogenous variables, i.e. $\mathbf{I}\subset\mathbf{X}$. The intervention on target $\mathbf{I}$ is given by $\mathbf{F}_{x}=\{f_{i}|X_{i}\notin\mathbf{I}\}\cup\{f_{j}|X_{j}\in\mathbf{I}\}.$ The intervention on variable $X_{j}$ is implemented with the change of the data generation mechanism from $f_{i}$ to a constant scalar. There are generally two types of interventions: hard and soft intervention. The hard intervention on target $\mathbf{I}$ is given by $\mathbf{F}_{x}=\{f_{i}|X_{i}\notin\mathbf{I}\}\cup\{f_{j}:=\alpha_{j}|X_{j}\in\mathbf{I}\}$. Here $\mathbf{\alpha}$ is the realization of variables in $\mathbf{I}$. The do-operation [19] $do(\mathbf{X}:=\bm{\alpha})$ represents the hard intervention that changes the SCM from $\mathcal{M}$ to $\mathcal{M}_{\mathbf{\alpha}}$. A hard intervention on a node breaks its causal connection with its parents.
@@
where $A_{i}$ is the continuous action variable for $X_{i}$ if $X_{i}\in\mathbf{I}$. Action $A_{i}$ with the corresponding value range $\mathcal{A}_{i}$ controls the change of the data generation mechanism of node $X_{i}$.
@@ 2.3 Function Network Bayesian Optimization
Similar to CBO, the structure of the functions in Function Network BO (FNBO, [4; 3]) is assumed to be given, and the parameters and specific form of each function are unknown. FNBO operates soft interventions and uses an expected improvement (EI) acquisition function to select actions. FNBO assumes that the system is noiseless, which could be a restrictive assumption in practice. Both FNBO and CBO belong to a broad research direction using intermediate observations from the computation of the unknown functions to improve the sample efficiency of vanilla BO algorithms [5].
@@ 3 Problem Statement
This section gives the problem statement of the proposed EXCBO algorithm. Compared to existing methods, our proposed CBO framework studies more general causal models as given in (3). Following existing CBO algorithms [2; 1; 23], the DAG $\mathcal{G}$ is assumed to be known. Similar to MCBO, our framework adopts GP surrogate models to guide the optimization of soft interventions controlled by an action vector $\mathbf{a}=\{a_{i}\}_{i=0}^{d}$ in reaching the optimal reward value. This section gives details about the problem settings studied in this paper.
3.1 Regularity Assumption
In our problem, the edges in SCM $\mathcal{M}$ is a fixed but unknown set of functions $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$ with $f_{i}:\mathcal{Q}\to\mathcal{X}_{i}$. We assume the structure of the SCM is known and the problem is causal sufficient, which means that there are no unobserved variables or confounders. We consider standard smoothness assumptions for the unknown functions $f_{i}$ defining over a compact domain $\mathcal{Q}$ [21]. For each node $i\in[d]$, we assume that $f_{i}()$ belongs to a reproducing kernel Hilbert space (RKHS) $\mathcal{H}_{k}$, a space of smooth functions defined on $\mathcal{Q}=\mathcal{Z}_{i}\times\mathcal{A}_{i}\times\mathcal{U}_{i}$. This means that $f_{i}\in\mathcal{H}_{k_{i}}$ is induced by a kernel function $k_{i}:\mathcal{Q}\times\mathcal{Q}\to\mathbb{R}$. We also assume that $k_{i}(s,s^{\prime})\leq 1$ for every $s,s^{\prime}\in\mathcal{Q}$. The RKHS norm of $f_{i}()$ is assumed to be bounded $\|f_{i}\|_{k_{i}}\leq\mathcal{B}_{i}$ for some fixed constant $\mathcal{B}_{i}>0$ [23]. We further assume that $\mathcal{A}_{i}$ and $\mathcal{U}_{i},i\in[d]$ are compact and bounded. The intervention target set $\mathbf{I}$ is assumed to a subset of $\mathbf{X}$, i.e., $|\mathbf{I}|\leq d-1$.
@@ 3.2 CBO via Exogenous Distribution Learning
Different from ANMs in existing CBO algorithms, such as MCBO [23] briefly reviewed in Section 5, we here propose to flexibly model the mapping $f_{i}()$ with introduced exogenous variables, for which we propose a new EXCBO framework for CBO through exogenous distribution learning. At step $t$ of CBO, the relation between actions $a_{i,t}$, $u_{i,t}$ and the observation $x_{i,t}$ is given by
$x_{i,t}=f_{i}(\mathbf{z}_{i,t},a_{i,t},u_{i,t})\ \ \forall i\in[d].$ (5)
Our surrogate model in EXCBO takes the exogenous variable $U_{i}$ as the input as shown in (5) and Figure 1. In our SCMs, the exogenous variables are not simply noise in the system. We assume they are important background or environmental factors that affect the output of the SCM. Additionally, each $U_{i}$ can be Gaussian or non-Gaussian and has multiple modes.
Using an example in biology, $x_{i,t}$ is the community size of a species of bacteria at time $t$, $u_{i,t}$ is the amount of a chemical compound in the environment that affects the growth of the bacteria but independent with other factors and unknown to us, and $a_{i,t}$ is a controllable factor, e.g. environment humidity.
Let $\mathcal{R}$ represent the root node set. As the parent of root nodes is null, we set $\mathbf{z}_{i,t}=\mathbf{0}$, if $i\in\mathcal{R}$. Similarly, we have $a_{d,t}=0$ at the target node $d$,
@@
Given an action vector $\mathbf{a}=\{a_{i}\}_{i=0}^{d}$ and $u=\{u_{i}\}_{i=0}^{d}$, the reward or target value is simply denoted with $y=\mathbf{F}(a,u)$. The problem becomes searching for an action vector maximizing the objective:
@@
where the expectation is taken over $u$. The goal of our method is to design a sequence of interventions $\{\mathbf{a}_{t}\}_{t=0}^{T}$ that achieves a high average expected reward. The learning of exogenous variable distribution provides a more accurate surrogate model given an SCM and observational data.
@@
3.3 Motivations for Exogenous Distribution Learning
In existing CBO models, exogenous distributions are ignored or marginalized to achieve the simplicity of interventions [2; 1; 23]. As mentioned in previous sections, we propose an encoder-decoder scheme to recover the exogenous variable of each endogenous node in an SCM. The distribution of the exogenous variable ($U$) is approximated with the density of the recovered surrogate exogenous variable ($\widehat{U}$) which is modeled with a flexible model, e.g. Gaussian Mixture. The recovered exogenous distribution improves the surrogate model in the approximation of the SCM. Thus different from existing methods [2; 1; 23] that typically rely on the ANM assumption, our EXCBO allows us to extend CBO to a broader scope of causal models beyond ANMs. In addition, with a more accurate surrogate model, our method could achieve improved reward and objective values.
@@ 4 Exogenous Distribution Learning
Given the observations of an endogenous node and its parents in a SCM, our method is to recover the distribution of the node’s exogenous variable. The learning of the exogenous distribution is implemented with Gaussian Processes. We first focus on the exogenous distribution learning regarding one node.
@@ 4.1 Exogenous Variable Recovery for One Node
According to (4), an endogenous variable $X_{i}$ either has an action variable $A_{i}$ or not. To better illustrate the idea with compact notations, we use $\mathbf{Z}_{i}$ to represent variable $X_{i}$’s parents and its action variable in this section (Section 4), i.e., $\mathbf{Z}_{i}=(\mathbf{Z}_{i},A_{i})$ if $X_{i}\in\mathbf{I}$. Hence the learning of node $X$’s exogenous distortion becomes a problem of recovering $U$’s distribution given the observations of $X$ and $\mathbf{Z}$ regarding equation $X=f(\mathbf{Z},U)$. To make our presentation concise, we have the following definition for the causal mechanism of the triplets $(\mathbf{Z},U,X)$ that appear as a node in an SCM.
@@
In a $\tau$-SCM, $\mathbf{Z}$ is the parent of $X$, and it may be multi-dimensional; $U$ is the exogenous variable of $X$. Different from the bijective generation mechanisms (BGM, [16]), $f()$ is not confined to be monotonic and invertible regarding $X$ and $U$ given a value of $\mathbf{Z}$.
The encoder-decoder framework (shown in Figure 2) is used to recover a surrogate of the exogenous variable.
@@
(EDS) Let $(\mathbf{Z},U,X,f)$ be a $\tau$-SCM. Let $\phi()$ be a regression model, $\phi():\mathcal{\mathbf{Z}}\to\mathcal{X}$, with mean $\mu_{\phi}()$ and variance $\sigma_{\phi}()$. We define $(\widehat{U},\phi,h,g)$ as encoder-decoder surrogate (EDS) model of the exogenous variable $U$ with the encoder $h():\mathcal{Z}\times\mathcal{X}\to\widehat{\mathcal{U}}$, $\widehat{U}:=h(\mathbf{Z},X):=\frac{X-\mu_{\phi}(\mathbf{Z})}{\sigma_{\phi}(\mathbf{Z})}$; and the decoder $g():\mathcal{Z}\times\widehat{\mathcal{U}}\to\mathcal{X}$, $X:=g(\mathbf{Z},\widehat{U})$.
With the observations of $X$ and its parent $\mathbf{Z}$, our method learns the encoder $h()$ that aims to recover the value of $U$, i.e., $\widehat{u}=h(\mathbf{z},x)$; our algorithm also learns the decoder $x=g(\mathbf{z},\widehat{u})$ that provides us a surrogate model of the mechanism $f()$.
Theorem 4.1 shows that the surrogate values of the exogenous variable $U$ can be recovered from the observations of node $X$ under the decomposable generation mechanism (DGM) assumption of $f$, i.e. $X=f(\mathbf{Z},U)=f_{a}(\mathbf{Z})+f_{b}(\mathbf{Z})f_{c}(U)$.
@@
Remark 4.1.
We use the distribution of $\widehat{U}=\mathbf{s}(U)=h(\mathbf{Z},X)$, i.e. $p(\widehat{U})$, to represent $p(U)$ in the surrogate model. With the decomposable assumption on $f()$, a $\tau$-SCM is counterfactually identifiable.
DGM is significantly different from the ANM [11] mechanism given by $X=f(\mathbf{pa}(X))+U$, which is a linear monotonic function of $U$ and widely used by many existing BO and CBO methods. In our DGM, $f_{c}()$ can be any continuous function, including nonlinear, non-convex, and non-monotonic function classes regarding $U$. Moreover, $f_{b}(\mathbf{Z})$ can be any continuous function as long as $f_{b}(\mathbf{Z})\neq 0$. The last term of a DGM, i.e. $f_{b}(\mathbf{Z})f_{c}(U)$, indicates that given a fixed value of the parent $Z$, the generated variable $X$’s variance is not only determined by $U$ but also by the value of $\mathbf{Z}$. Therefore, with $f_{b}(\mathbf{Z})$ DGM represents a larger class of mechanisms that both $U$ and $\mathbf{Z}$ affect the variance of the generated variable $X$. These features substantially extend the identification of exogenous variable $U$ to much broader nonlinear and non-monotonic data-generation mechanism classes compared to ANM (linear) and BGM (monotonic). Figure 3 gives the relationship between different mechanisms.
@@
In our surrogate model, $f()$ is approximated with the learned $g()$ and $\widehat{u}$, i.e., $x=f(\mathbf{z},u)=g(\mathbf{z},\widehat{u})=g(\mathbf{z},\mathbf{s}(u))$. With the monotonic assumption on $f()$, we can extend the EDS framework to BGMs by following the analysis in [14; 16; 17; 8].
Theorem 4.2.
Let $(\mathbf{Z},U,X,f)$ be a $\tau$-SCM. $\forall\mathbf{z}\in\mathcal{Z}$, $f(\mathbf{z},\cdot)$ is differentiable and strictly monotonic regarding $u\in\mathcal{U}$. Let $(\widehat{U},\phi,h,g)$ be an EDS surrogate of $U$. We further assume that $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$. Then $\widehat{U}=h(\mathbf{Z},X)$ is a function of $U$, i.e., $\widehat{U}=\mathbf{s}(U)$, and $\mathbf{s}()$ is an invertible function.
The proofs of Theorem 4.1 and Theorem 4.2 can be found in the Appendix. The proposed exogenous variable surrogate $\widehat{U}$ and the encoder $h()$ work under either the DGM conditions in Theorem 4.1 or the BGM [16] conditions in Theorem 4.2. This means that the proposed EDS model is suitable for ANMs, DGMs, and BGMs. For a BGM model, we need to enforce independence between $\widehat{U}$ and $\mathbf{Z}$, i.e. $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$, to recover $U$ according to Theorem 4.2. It could be achieved by adding an independence regularization on both variables, which may introduce additional computation costs.
We use EDS to represent the $\tau$-SCMs that are counterfactually identifiable by using EDS with or without the $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$ condition and Figure 3 gives the scopes of different mechanisms including EDS. If $f()$ does not meet the DGM and the BGM assumptions, the learned $\widehat{U}=\frac{X-\mu_{\phi}(\mathbf{Z})}{\sigma_{\phi}(\mathbf{Z})}$ may not be independent of the parent variable $\mathbf{Z}$, and we may not get an accurate surrogate model for the SCM. It may decrease the possibility of attaining the optimal objective value of $y$ using limited training samples.

### v10

@@ Abstract
Maximizing a target variable as an operational objective in a structural causal model is an important problem. Causal Bayesian Optimization (CBO) methods either rely on interventions that alter the causal structure to maximize the reward; or introduce action nodes to endogenous variables so that the data generation mechanisms are adjusted to achieve the objective. This paper introduces a novel method to learn the distribution of exogenous variables, which is typically marginalized through expectation or ignored by existing CBO methods. Exogenous distribution learning improves the approximation accuracy of structural causal models in a surrogate model that is usually trained with limited observational data. Moreover, the learned exogenous distribution extends existing CBO to general causal schemes beyond simple Additive Noise Models (ANMs). The recovery of exogenous variables allows us to use a more flexible prior for noise or unobserved hidden variables. We develop a new CBO method by leveraging the learned exogenous distribution. Experiments on different datasets and applications show the benefits of our proposed method.
@@ 1 Introduction
Bayesian Optimization (BO) has broad applications, such as automating modern industrial processes, drug discovery, and more general synthetic biology, that require optimizing black-box objective functions [15; 3; 10]. In many real-world problems, we have some structural knowledge of the unknown functions that can be leveraged to improve BO efficiency. CBO has been developed to utilize the structural information of the objective [1; 2; 22]. CBO combines ideas from causal inference, uncertainty quantification, and sequential decision-making. It generalizes Bayesian Optimization, which treats the input variables of the objective function as independent, to scenarios where causal information between input variables is available [1]. CBO has been developed for optimizing medical and ecological interventions as illustrative examples [1; 2].
@@ 1.1 Approach and Contributions
In this paper, we develop an EXogenous distribution learning augmented Causal Bayesian Optimization method, EXCBO. Given observation data points of a structural causal model (SCM, [18; 17]), the exogenous variable associated with each endogenous node is recovered with an encoder-decoder framework. The learned distribution of the exogenous variable is then approximated with the density of the recovered exogenous variable that is parameterized with a flexible model, e.g. Gaussian Mixture. The recovered exogenous distribution improves the surrogate model’s accuracy in the SCM approximation. Different from existing CBO methods [1; 2; 22] that typically focus on Additive Noise Models (ANMs, [11]), our method enables us to extend CBO to general causal models beyond ANMs. Our flexible surrogate model has the benefit of better causal inference regarding CBO updating.
@@
  • The flexibility of learning exogenous distributions naturally extends the proposed CBO framework to general causal models beyond ANMs.
@@
In the following parts of this paper, Section 2 presents the background and related work; Section 3 sets up the problem studied in this paper with the introduction of the proposed CBO framework; Section 4 presents the exogenous recovery method; The proposed SCM algorithm, EXCBO, is detailed in Section 5 with theoretical analysis of EXCBO given in Section 6; Section 7 presents experimental results; Section 8 lists additional remarks; Section 9 concludes the paper.
2 Background
@@ 2.1 Structural Causal Model
An SCM is denoted by $\mathcal{M}=(\mathcal{G},\mathbf{F},\mathbf{V},\mathbf{U})$. Here $\mathcal{G}$ is a directed acyclic graph (DAG), and $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$ stands for $d+1$ mechanisms of all variables. $\mathbf{V}$ is the endogenous variable set, and $\mathbf{U}$ is the exogenous (background) variable set. Samples of the $i$th endogenous variable are generated with
@@
Here $[d]=\{0,1,...,d\}$. We use $X_{i}$ to represent the $i$th variable as well as the $i$th node in $\mathcal{G}$, and $\mathbf{pa}(i)$ stands for the parents of $i$, and $\mathbf{ch}(i)$ is the children of node $X_{i}$ regarding $\mathcal{G}$. In addition, we have $U_{i}\perp\!\!\!\perp X_{i}$ and $U_{i}\perp\!\!\!\perp U_{j},\forall i\neq j$. $f_{i}$ is a mapping from $\mathbb{R}^{|\mathbf{pa}(i)|+1}\rightarrow\mathbb{R}$. $\mathcal{X}_{i}$, $\mathbf{\mathcal{Z}}_{i}$, and $\mathcal{U}_{i}$ stand for the value domain of $X_{i}$, $\mathbf{Z}_{i}$, and $U_{i}$, respectively. We further assume that the expectation of $X_{i},i\in[d]$, i.e. $\mathbb{E}[X_{i}]$, exists. Existing CBO methods [1; 22] usually assume that the exogenous variable is an ANM [11] to the data generation mechanism, i.e., $X_{i}=f_{i}(\mathbf{Z}_{i})+U_{i},\ \text{with}\ U_{i}\sim\mathcal{N}(0,1).$
@@ 2.2 Intervention
In SCM $\mathcal{M}$, let $\mathbf{I}$ be a subset of endogenous variables, i.e. $\mathbf{I}\subset\mathbf{U}$. The intervention on target $\mathbf{I}$ is given by $\mathbf{F}_{x}=\{f_{i}|X_{i}\notin\mathbf{I}\}\cup\{f_{j}|X_{j}\in\mathbf{I}\}.$ The intervention on variable $X_{j}$ is implemented with the change of the data generation mechanism from $f_{i}$ to a constant scalar. There are generally two types of interventions: hard and soft intervention. The hard intervention on target $\mathbf{I}$ is given by $\mathbf{F}_{x}=\{f_{i}|X_{i}\notin\mathbf{I}\}\cup\{f_{j}:=\alpha_{j}|X_{j}\in\mathbf{I}\}$. Here $\mathbf{\alpha}$ is the realization of variables in $\mathbf{I}$. The do-operation [18] $do(\mathbf{X}_{\mathbf{I}}:=\bm{\alpha})$ represents the hard intervention that changes the SCM from $\mathcal{M}$ to $\mathcal{M}_{\mathbf{\alpha}}$. A hard intervention on a node breaks its causal connection with its parents.
@@
where $A_{i}$ is the continuous action variable for $X_{i}$ if $X_{i}\in\mathbf{I}$. Action $A_{i}$ with the corresponding value domain $\mathcal{A}_{i}$ controls the change of the data generation mechanism of node $X_{i}$. The soft intervention can be represented with Pearl’s do-operation [18] $do\big{(}\mathbf{X}_{\mathbf{I}}:=\mathbf{f}(\mathbf{Z}_{\mathbf{I}},A,U_{\mathbf{I}})\big{)}$
@@ 2.3 Function Network Bayesian Optimization
Similar to CBO, the structure of the functions in Function Network BO (FNBO, [4; 3]) is assumed to be given, and the parameters and specific form of each function are unknown. FNBO operates soft interventions and uses an expected improvement (EI) acquisition function to select actions. FNBO assumes that the system is noiseless, which could be a restrictive assumption in practice. Both FNBO and CBO belong to a broad research direction using observations of a function network to improve the sample efficiency of vanilla BO algorithms [5].
@@ 3 Problem Statement
Following existing CBO algorithms [1; 2; 22], the DAG $\mathcal{G}$ is assumed to be known in this paper. Similar to MCBO [22], our framework adopts GP surrogate models to guide the optimization of soft interventions controlled by an action vector $\mathbf{a}=\{a_{i}\}_{i=0}^{d}$ in reaching the optimal reward value. This section gives details about the problem settings studied in this paper.
3.1 Motivations for Exogenous Distribution Learning
In existing CBO models, exogenous distributions are ignored or marginalized to achieve the simplicity of interventions [1; 2; 22]. As mentioned in previous sections, we propose an encoder-decoder scheme to recover the exogenous variable of each endogenous node in an SCM. The distribution of an exogenous variable ($U_{i}$) is approximated with the density of the recovered surrogate exogenous variable ($\widehat{U}_{i}$) which is modeled with a flexible model, e.g. Gaussian Mixture. The recovered exogenous distribution improves the surrogate model in the approximation of the SCM. Thus different from existing methods [1; 2; 22] that typically rely on the ANM assumption, our EXCBO allows us to extend CBO to a broader scope of causal models beyond ANMs. In addition, with a more accurate surrogate model, our method could achieve improved reward and objective values.
3.2 Decomposable Generation Mechanism
In our problem, we assume the DAG structure of the SCM $\mathcal{M}$, i.e. $\mathcal{G}$, is given, and the edges or the mechanisms of the endogenous variables, i.e. $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$, are fixed but unknown. We further assume that the problem is causal sufficient, which means that there are no unobserved endogenous variables or confounders. We present the definition of decomposable generation mechanism (DGM) used in this paper as follows.
Definition 3.1.
(DGM) The data generation function $f$ follows decomposable generation mechanism (DGM) if $X=f(\mathbf{Z},U)=f_{a}(\mathbf{Z})+f_{b}(\mathbf{Z})f_{c}(U)$, and the mappings are $f_{a}():\mathcal{Z}\to\mathbb{R}$, $f_{b}():\mathcal{Z}\to\mathbb{R}$, and $f_{c}():\mathcal{U}\to\mathbb{R}$.
In a DGM, $f_{c}()$ can be any continuous function, including nonlinear and non-monotonic one-dimensional function classes regarding $U$. $f_{b}(\mathbf{Z})$ can be any continuous function with $f_{b}(\mathbf{Z})\neq 0$. The last term of a DGM, i.e. $f_{b}(\mathbf{Z})f_{c}(U)$, indicates that given a fixed value of the parent $Z$, the generated variable $X$’s variance is not only determined by $U$ but also by the value of $\mathbf{Z}$. Therefore, with $f_{b}(\mathbf{Z})$ DGM represents a large class of mechanisms that both $U$ and $\mathbf{Z}$ affect the variance of the generated variable $X$. DGM is different from the LSNMs (location-scale noise models, [12]), which adopts a locally linear relationship between the exogenous variable $U$ and the endogenous variable $X$.
We show that the distributions of exogenous variables can be recovered if the data generation mechanism $f$ in each node (3) belongs to DGM in Section 4.1.
@@ 3.3 CBO via Exogenous Distribution Learning
Unlike ANMs in existing CBO algorithms, e.g. MCBO [22] (briefly reviewed in Section 2.4), we here propose to flexibly model the mapping $f_{i}()$ with exogenous variables, for which we propose a new EXCBO framework for CBO through exogenous distribution learning.
At step $t$ of CBO, the relation between actions $\mathbf{a}_{i,t}$, $\mathbf{u}_{i,t}$ and the observation $\mathbf{x}_{i,t}$ is given by
$\mathbf{x}_{i,t}=\mathbf{f}_{i}(\mathbf{z}_{i,t},\mathbf{a}_{i,t},\mathbf{u}_{i,t})\ \ \forall i\in[d].$ (5)
Our surrogate model in EXCBO takes the exogenous variable $U_{i}$ as the input as shown in (5) and Figure 1. Let $\mathcal{R}$ represent the set of the root nodes. As the parent of root nodes is null, we set $\mathbf{z}_{i,t}=\mathbf{0}$, if $i\in\mathcal{R}$. Similarly, we have $a_{d,t}=0$ at the target node $d$,
@@
Given an action vector $\mathbf{a}=\{a_{i}\}_{i=0}^{d}$ and $\mathbf{u}=\{u_{i}\}_{i=0}^{d}$, the reward or target value is simply denoted with $y=\mathbf{F}(\mathbf{a},\mathbf{u})$. The problem becomes searching for an action vector maximizing the objective:
@@
where the expectation is taken over $u$. The goal of our method is to design a sequence of interventions $\{\mathbf{a}_{t}\}_{t=0}^{T}$ that achieves a high average expected reward. In Section 4, we will show that if a mechanism ${f_{i}}$ in $\mathbf{F}=\{f_{i}\}_{i=0}^{d}$ meets the DGM assumption, the exogenous variable $U_{i}$ and its distribution can be learned from observations. With the learned distributions of all the exogenous variables, each action’s reward can be evaluated in the operations of equation (6).
@@ 4 Exogenous Distribution Learning
Given the observations of an endogenous node and its parents in a SCM, our method is to recover the distribution of the node’s exogenous variable. The learning of the exogenous distribution is implemented with GPs. We first focus on the exogenous distribution learning regarding one node.
@@ 4.1 Exogenous Variable Recovery for One Node
According to (4), an endogenous variable $X_{i}$ either has an action variable $A_{i}$ or not. To better illustrate the idea with compact notations, we use $\mathbf{Z}_{i}$ to represent variable $X_{i}$’s parents and its action variable in this section (Section 4), i.e., $\mathbf{Z}_{i}=(\mathbf{Z}_{i},A_{i})$ if $X_{i}\in\mathbf{I}$. Hence the learning of node $X$’s exogenous distortion becomes a problem of recovering $U$’s distribution given the observations of $X$ and $\mathbf{Z}$ regarding equation $X=f(\mathbf{Z},U)$. To make our presentation concise, we have the following definition for the causal mechanism of the triplets $(\mathbf{Z},U,X)$ that appears as a node in an SCM.
@@
In a $\tau$-SCM, $\mathbf{Z}$ is the parent of $X$, and it may be multi-dimensional; $U$ is the exogenous variable of $X$, and it is different from the bijective generation mechanisms (BGM, [16]). In BGM with data generation mechanism $X=f(\mathbf{Z},U)$, $f()$ is required to be monotonic and invertible regarding $X$ and $U$ given a value of $\mathbf{Z}$. The encoder-decoder framework (shown in Figure 2) is used to recover a surrogate of the exogenous variable.
@@
(EDS) Let $(\mathbf{Z},U,X,f)$ be a $\tau$-SCM. Let $\phi()$ be a regression model, $\phi():\mathbf{\mathcal{Z}}\to\mathcal{X}$. We assume that $\mathbb{E}[X]$ exists and $\phi()$ has sufficient expressiveness to capture the regression mean $\mu_{\phi}()$ and variance $\sigma_{\phi}()$. We define $(\widehat{U},\phi,h,g)$ as encoder-decoder surrogate (EDS) model of the exogenous variable $U$ with the encoder $h():\mathcal{Z}\times\mathcal{X}\to\widehat{\mathcal{U}}$, $\widehat{U}:=h(\mathbf{Z},X):=\frac{X-\mu_{\phi}(\mathbf{Z})}{\sigma_{\phi}(\mathbf{Z})}$; and the decoder $g():\mathcal{Z}\times\widehat{\mathcal{U}}\to\mathcal{X}$, $X:=g(\mathbf{Z},\widehat{U})$.
With the observations of $X$ and its parent $\mathbf{Z}$, our method learns the encoder $h()$ that aims to recover the value of $U$, i.e., $\widehat{u}=h(\mathbf{z},x)$; our algorithm also learns the decoder $x=g(\mathbf{z},\widehat{u})$ that provides us a surrogate model of the mechanism $f()$. Theorem 4.1 shows that the surrogate values of the exogenous variable $U$ can be recovered from the observations of node $X$ under the DGM assumption of $f$.
@@
We use the distribution of $\widehat{U}=\mathbf{s}(U)=h(\mathbf{Z},X)$, i.e. $p(\widehat{U})$, to represent $p(U)$ in the surrogate model. In our surrogate model, $f()$ is approximated with the learned $g()$ and $\widehat{u}$, i.e., $x=f(\mathbf{z},u)=g(\mathbf{z},\widehat{u})=g(\mathbf{z},\mathbf{s}(u))$. Figure 3 gives the relationship between different mechanisms. These features substantially extend the identification of exogenous variable $U$ to much broader nonlinear and non-monotonic data-generation mechanism classes compared to ANM (linear) and BGM (monotonic, [16]). DGM is significantly different from the ANM [11] mechanism given by $X=f(\mathbf{pa}(X))+U$, which is a linear monotonic function of $U$ and widely used by many existing BO and CBO methods. We use EDS to represent the $\tau$-SCMs with the exogenous distribution that can be approximately recovered by using EDS (with or without the condition $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$) and Figure 3 gives the scopes of different mechanisms including EDS.